OpenStudy (anonymous):

Is it possible for x=-2 to be in the domains of the function: R(x)= 3x^3-5x+6 -------------- x+2

OpenStudy (anonymous):

Only if 3x^3-5x+6 has a factor of x+2

OpenStudy (anonymous):

But if x + 2 was a factor, wouldn't that cause a hole, meaning its undefined..

OpenStudy (anonymous):

Good point. It's not a factor anyway so... :)

OpenStudy (anonymous):

I was thinking in terms of limits.... =/

OpenStudy (anonymous):

So the answer is no because why?

OpenStudy (anonymous):

Can you have a 0 in the denominator of a fraction?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:) That's why...

OpenStudy (anonymous):

The question just asks if its possbile

OpenStudy (anonymous):

oh Thanks sry i have never used this before.

OpenStudy (anonymous):

the second part of that ? is

OpenStudy (anonymous):

np... :) I appreciate that you did ask "why" though. Many people who post here don't care to know why....they just want the answer.

OpenStudy (anonymous):

second part is what?

OpenStudy (anonymous):

H(x)= square root x+1

OpenStudy (anonymous):

Yeah I want to learn how to do them not just get the answers.

OpenStudy (anonymous):

I have a few more but cant figure ourt how to type them to you. lol

OpenStudy (anonymous):

is that sqrt(x + 1) or sqrt(x) + 1? ANd what are you looking for to be in the domain?

OpenStudy (anonymous):

it just says x+1 with no ()s

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Is the x+1 under the sqaure root sign, or is the 1 on the outside?

OpenStudy (anonymous):

OH! Sorry all under the square root sign

OpenStudy (anonymous):

we use the parentheses to show that... (for future reference when typing these) :) So you have sqrt(x + 1). Can you take the square root of a negative number?

OpenStudy (anonymous):

I didnt think so.

OpenStudy (anonymous):

you cant have a neg inside the sqrt

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

In order for a number to be part of the domain, you have to be able to evaluate the expression at that number. So for x = -2 to be in the domain, you have to be able to evaluate sqrt(-2 + 1) = sqrt (-1).

OpenStudy (anonymous):

Right...so since you can't take the sqrt of a negative...is x = -2 in the domain? ;)

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

lol...try again with confidence!

OpenStudy (anonymous):

I have to write out why or why not. So I would just answer this by saying you cant take the sqrt of a neg #?

OpenStudy (anonymous):

Haha. No!

OpenStudy (anonymous):

That's beter! we'll make a mathematician out of you yet!

OpenStudy (anonymous):

What do you do for a living cause you should be a teacher. lol I like the way you teach!

OpenStudy (anonymous):

I used to be a teacher...college math. then i decided to join the military for a change. when i get out, i plan to go back to teaching. :)

OpenStudy (anonymous):

Ok here is another one.. Suppose we are given two functions f(x)= 3x+2-4/x and g(x)= 2-3x. find the indicated values: f(a+b)

OpenStudy (anonymous):

if i asked you to find f(3), could you?

OpenStudy (anonymous):

Well you are VERY good at it and I like that you throw in some positiveness!! ha.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then do the same thign...only use a+b instead of 3 :)

OpenStudy (anonymous):

I am not very good with fractions :(

OpenStudy (anonymous):

ohh. I get it! Der.

OpenStudy (anonymous):

embrace the fraction!

OpenStudy (anonymous):

I am kinda confused on the last part thou

OpenStudy (anonymous):

4/a+b? <-- is that how I would write it?.

OpenStudy (anonymous):

is the problem \[3x+2-\frac{4}{x}\]? Then the answer would be \[3a+3b+2-\frac{4}{a+b}\]

OpenStudy (anonymous):

hey how did you do that? lol

OpenStudy (anonymous):

lol...the equation editor :)

OpenStudy (anonymous):

I cant fail this assignment so thank you so much for helping me!!

OpenStudy (anonymous):

\frac{}{} is the code for a fraction. put the numerator in the first set of curlies and the denominator in the second set

OpenStudy (anonymous):

wrap it inside of a \[

OpenStudy (anonymous):

and a \]

OpenStudy (anonymous):

so... slash[\frac{3}{4}\] (replacing the wotrd "slash" with a "\") will give you \[\frac{3}{4}\]

OpenStudy (anonymous):

The second part of that ? is: g(a)-g(a-b)+f(a

OpenStudy (anonymous):

Go slowly...evaluate each part [g(a), g(a-b) and f(a)] then connect them together as requested...

OpenStudy (anonymous):

Wheter or not you need to combine the fractions is up to your instructor. personally, unless i'm working with common denominators, i wouldn't have my students worry about it.

OpenStudy (anonymous):

this one really confuses me

OpenStudy (anonymous):

the purpose of the problem is to get you to understand the mechanics of evaluating at non-numbers...not to cause problems with nasty denominators.

OpenStudy (anonymous):

what is g(a)? we'll start there

OpenStudy (anonymous):

I know i need to plug that in where x is but do i plug that whole thing in

OpenStudy (anonymous):

for g(a), you just plug a in for x...

OpenStudy (anonymous):

so g(a)=-3(a)

OpenStudy (anonymous):

g(x) = 2-3x...

OpenStudy (anonymous):

opps . i left out the 2

OpenStudy (anonymous):

so in place of the x i put a. g(a)=2-3(a)?

OpenStudy (anonymous):

right...so what is g(a-b)?

OpenStudy (anonymous):

g(a-b)=2-3(a-b)?

OpenStudy (anonymous):

distribute...

OpenStudy (anonymous):

ga-gb=2-3a+3b?

OpenStudy (anonymous):

well...you don't distribute the g(a-b) part...since g is a function...ut the right side is correct.

OpenStudy (anonymous):

finally, what is f(a)?

OpenStudy (anonymous):

Ohh. Opps. Im not good at math at all. Splendid in english thou :)

OpenStudy (anonymous):

f(a)=2-3a?

OpenStudy (anonymous):

we'll get you there. :)

OpenStudy (anonymous):

nope...that is the function g...not the function f

OpenStudy (anonymous):

remember, the function f is f(x) = 3x+2-(4/x)

OpenStudy (anonymous):

Uh. Help?!

OpenStudy (anonymous):

lol...g(x) and f(x) are two different functions. just like your cofee maker and dishwasher have two different functions.

OpenStudy (anonymous):

lol...g(x) and f(x) are two different functions. just like your cofee maker and dishwasher have two different functions.

OpenStudy (anonymous):

lol...g(x) and f(x) are two different functions. just like your cofee maker and dishwasher have two different functions.

OpenStudy (anonymous):

lol...g(x) and f(x) are two different functions. just like your cofee maker and dishwasher have two different functions.

OpenStudy (anonymous):

haha um

OpenStudy (anonymous):

i don't know why that happened.

OpenStudy (anonymous):

g(x) = 2-3x and f(x) = 3x+2-(4/x), right?

OpenStudy (anonymous):

I didnt know the two functions went together

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thats tricky.

OpenStudy (anonymous):

putting them together is like making coffee then putting the pot in the dishwasher. so what we do is we evauate each piece...then string them together.

OpenStudy (anonymous):

Given the functions f(x) = 3x + 2- 4/x and g(x) = 2 - 3x, fnd the indicated values. a) f(a + b) b) g(a) - g(a - b) + f(a)

OpenStudy (anonymous):

right...so for part b) so far we have that...

OpenStudy (anonymous):

a) is the first one and b) is the second they mushed together

OpenStudy (anonymous):

g(a) = 2-3a

OpenStudy (anonymous):

g(a-b) = 2-3a+3b

OpenStudy (anonymous):

what is f(a)? next we'll mash them together...